1. An airplane flying at 22,000 feet descended 8,500 feet and then ascended 5,000 feet. What is the new altitude of the airplane?
Answer by Boreal(15235)

22000-8500=13,500
13500+500=18,500
or
add the plus ones 22000+5000=27000
subtract the minus one -8500
18,500 feet.